suppose X=3 with probabiltiy 0.6 and X=9 with probability 0.7. obtain E[x] and Var[X]

probabilriy .6 and .4
Answer #1

E(x) =xP(x) =3*0.6+9*0.4 =5.4

E(x2)=x2P(x) =3^2*0.6+9^2*0.4 =37.8

Var(x)=E(x2)-(E(x))2 =37.8-5.4^2 =8.64